OpenStudy (anonymous):

How many different 4-digit sequences can be formed using the digits 0, 1,..., 6 if repetition of digits is allowed? a. 1296 c. 24 b. 20 d. 2401

OpenStudy (anonymous):

First, identify how many different "choices" you have for each digit of the 4 digit number.

OpenStudy (anonymous):

its 7x7x7x7

OpenStudy (anonymous):

2041

OpenStudy (anonymous):

There ya go.

OpenStudy (anonymous):

thanks for you help :)

OpenStudy (anonymous):

My pleasure :)

OpenStudy (anonymous):

A card is picked at random from a standard deck of cards. What is the sample space for this experiment if you were trying to find P(face card)?

OpenStudy (anonymous):

@Ashleyisakitty @austinL @adrynicoleb @annas @Abdulhameed @bbcream14 @Biokiiller396

OpenStudy (anonymous):

@Compassionate @Vandreigan

OpenStudy (anonymous):

im sorry i dont no this one but i can get some people who can help you

OpenStudy (anonymous):

@Whitemonsterbunny17 @TheOcean @elementwielder

OpenStudy (anonymous):

I'm trying to remember the exact definition of sample space. I could calculate P(face), but I'm not certain of how to answer this, off the top of my head.

OpenStudy (anonymous):

I guess you'd define the sample space as {face card, not a face card}, since we're only interested in whether it's a face card or not.

OpenStudy (anonymous):

In a certain lottery, 5 numbers between 1 and 13 inclusive are drawn. These are the winning numbers. How many different selections are possible? Assume that the order in which the numbers are drawn is not important. a. 1287 c. 371,293 b. 154,440 d. 120

OpenStudy (anonymous):

what about this one?

OpenStudy (anonymous):

I'm assuming this is without replacement. Use the binomial coefficient: \[\left(\begin{matrix}N \\ k\end{matrix}\right) = \frac{N!}{k!(N-k)!}\] This gives you the number of possible ways to choose k different things from set of size N. In your case, you want to choose 5 numbers from a set of 13 numbers.